OpenStudy (anonymous):

cos^2x-sinxcosx-sin^2x

OpenStudy (anonymous):

Are you certain this is entirely accurate? (Speaking primarily about the second subtraction, before sin^2x)

OpenStudy (anonymous):

i dont think so that you stated it rightly..............

OpenStudy (anonymous):

cos^2x = sin^2x + sinxcosx 1= sin^2x/cos^2x + sinxcosx / cos^2x 1 = tan^2x + sinx / cosx 1= tan^2x + tanx jsut played with it ^^